OpenStudy (anonymous):

Bacteria triple every 2 hours, How long does it take the bacteria population to quadruple?

OpenStudy (anonymous):

Is there a starting population for the bacteria?

OpenStudy (anonymous):

Don't need starting population.

OpenStudy (anonymous):

dN/dt = k N dN/N = k t N = No exp(k t) N/No = exp(k t) N/No = 3 = exp( 2 k) 2 k = ln (3) = 1.10 k = 0.55 4 = exp (0.55 t) ln (4) = 0.55 t t = 2.52 Check this to make sure I got it right and you know how it's done.

OpenStudy (wolf1728):

Population = n * (sq root (3)^hours) After 2 hours population = n * 1.732^2 = 3

OpenStudy (wolf1728):

To quadruple, 4 = sqrt(3)^x log (4) = x * log(sqrt(3) x = log (4) / log (sqrt(3)) x = 0.6020599913 / 0.239 x = 2.52 hours

OpenStudy (anonymous):

We agree.

OpenStudy (wolf1728):

Yes, we got the same answer

OpenStudy (anonymous):

what's the sqrt for?

OpenStudy (anonymous):

Essentially, you worked with 2^(a t) and I worked with e^(k t), and these can be made equivalent.

OpenStudy (anonymous):

why do you have to sqrt 3?